# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

### v3

@@ Abstract
The evolution of artificial intelligence (AI) and neural network theories has revolutionized the way software is programmed, shifting from a hard-coded series of codes, Software 1.0, to a vast neural network, Software 2.0. However, this transition in engineering software has faced challenges such as data scarcity, multi-modality of data, low model accuracy, and slow inference. Here, we propose a new network based on interpolation theories and tensor decomposition, the interpolating neural network (INN) to open the new era of Engineering Software 2.0 that unifies training, solving, and calibration. Instead of interpolating training data, a common notion in computer science, INN interpolates grid points in the physical space whose coordinates and values are trainable. INN features orders of magnitude fewer trainable parameters (or degrees of freedom for solving), faster training/solving, less inference cost, smaller memory footprint, and higher model accuracy compared to multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Various numerical experiments that cover computer science and engineering domains demonstrate that INN can solve over Zetta scale $(10^{21})$ partial differential equations and train/calibrate a dataset with extraordinary accuracy but fewer parameters using only a single graphics processing unit (GPU).
@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies within the field of engineering and science introduces distinct challenges. These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6, 7], slow inference (function evaluation) time that challenges real-time control [8, 9, 10, 11] and statistical modeling, and resource-heavy model architecture [12, 13, 14, 15].
@@
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibration in science/engineering problems."
@@
As one promising enabler of Engineering Software 2.0, we introduce interpolating neural networks (INN). The key ideas behind INN are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message-passing operation as a form of classical interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes. Figure 1 shows the pictorial explanation of the approach. Note that one can also extrapolate for a certain range if the interpolation points are located outside of the range of training data. INN facilitates three major tasks in computational science and engineering: 1) train (or learn) on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solve high-dimensional partial differential equations (PDEs) with statistics involved, and 3) calibrate (or solve inverse problems) for uncertain parameters.
When the input domain is discretized with a regular mesh (see the special case of Figure 1), INN leverages the separation of variables and tensor decomposition (TD) [16, 17, 18, 19] to convert the growth rate of the computational cost of a high-dimensional problem from exponential to linear. This enables solving extremely challenging (over exascale, $10^{18}$) high-dimensional PDEs or training data with hundreds of input variables within the interpolation framework.
Our results demonstrate that INN can markedly enhance training/inference speed and model accuracy compared to the multilayer perceptron (MLP) or physics-informed neural network (PINN)[20] by dramatically reducing the number of trainable parameters and network connectivity. This also results in a huge reduction in the usage of random-access memory (RAM) and disk storage, facilitating the transferability of a trained/solved model across multiple workstations. Furthermore, INN can adapt its network structure further to improve model accuracy or incorporate statistical behavior. INN has a huge potential of replacing MLP and convolutional neural network (CNN) [21] in modern AI models with the aforementioned benefits.
@@
Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Networks (INN). The INN box illustrates the graph representation of the input domain discretized with an arbitrary Euclidean mesh (left) and a regular mesh as a special case (right). Node and edge features are given from the discretized mesh. After $Q$-step message passing, each node $j$ will store an interpolation function $N^{[Q](j)}(\bm{x})$. Finally, the readout operation sums the product of the interpolation functions and nodal values $\bm{u}^{(j)}$. The superscripts with square brackets [] and parentheses () denote the message passing step and graph node index, respectively. The interpolation operator $\mathcal{J}$ denotes that $\mathcal{J}\bm{u}(\bm{x})$ is a function that interpolates discrete values of $\bm{u}^{(j)}$. The INN trainer employs data-driven loss functions (e.g., mean squared error loss for regression) while the INN solver adopts a residual loss of a partial differential equation (PDE). A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
2 INN: Unifying Training/Learning, Solving, and Calibration
@@
Interpolating Neural Network (INN) generalizes numerical methods such as finite element method (FEM) [22, 23, 24] and neural networks such as multi-layer perception (MLP) as a graph network operation. Although the two (i.e., numerical methods and neural networks) have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1.
@@
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input, as illustrated in the INN box of Figure 1. Note that this conversion from mesh to graph has been widely accepted in literature [25, 26, 27]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$, which stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
@@
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [28]. In contrast, INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state. In other words, the message passing operation of INN is functional, an operation that returns a function. A general $Q$-step message passing of INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
@@
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information for visual illustrations). The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$. This operation can be any interpolation method such as Lagrange polynomials [29, 22], meshfree [30, 31, 32], and splines [33, 34]. The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [35, 36] and convolution hierarchical deep learning neural networks (C-HiDeNN) [7, 18]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 18]. See Supplementary Information for various choices of the message passing operation.
@@
$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L}.$ (2)
Here, the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$). The interpolation operator $\mathcal{J}$ will designate an interpolated field output throughout this paper. The readout operation can be written as a tensor contraction (or matrix multiplication):
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}^{T}\bm{\mathcal{X}(x)},$ (3)
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{J\times L},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. The $\bm{U}^{T}\in\mathbb{R}^{L\times J}$ refers to the matrix transpose. Here, $\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second order of tensor $\bm{U}^{T}$ and the first order of tensor $\bm{\mathcal{X}(x)}$, while $\bm{U}^{T}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication. We note that $\bm{U}$ is a stack of nodal values $\bm{u}^{j}$, while $\bm{\mathcal{X}(x)}$ is a function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to the r-adaptivity in FEM [23]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, and calibration.
@@
When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [16]. One of the widely used TD methods is the Tucker decomposition [37, 38]. Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values $\bm{u}$ will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
@@
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a full trainable matrix typically smaller than the original nodal discretization ($M_{1}<J_{1},M_{2}<J_{2},M_{3}<J_{3}$). This enables data compression. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is the 1-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{\bm{u}}^{(1)}&\prescript{1}{i}{\bm{u}}^{(2)}&\cdots&\prescript{1}{i}{\bm{u}}^{(J_{i})}\\ \prescript{2}{i}{\bm{u}}^{(1)}&\prescript{2}{i}{\bm{u}}^{(2)}&\cdots&\prescript{2}{i}{\bm{u}}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{\bm{u}}^{(1)}&\prescript{M_{i}}{i}{\bm{u}}^{(2)}&\cdots&\prescript{M_{i}}{i}{\bm{u}}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}\times 1}$. When there is more than one output $(L>1)$, the interpolated field becomes a vector of $L$ elements:
@@
$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}.$
@@
Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [39, 40, 41] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements. If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the CP decomposition yields:
$\mathcal{J}\bm{u}(\bm{x})$ $=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\sum^{M}_{m=1}{\prod^{I}_{i=1}{\mathcal{J}\prescript{m}{i}{u}(x_{i})}},$ (7)
an inner product of the three interpolated vectors of length $M$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is a one-dimensional function. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with multiple one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$.
Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output relationship. The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization. Assume that the nodal coordinates $(\bm{x}^{j})$ are fixed.

### v4

# Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration

@@ Abstract
Artificial intelligence (AI) has revolutionized software development, shifting from task-specific codes (Software 1.0) to neural network-based approaches (Software 2.0). However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations. We introduce an interpolating neural network (INN), grounded in interpolation theory and tensor decomposition, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration. INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Demonstrated in metal additive manufacturing, INN rapidly constructs an accurate surrogate model of Laser Powder Bed Fusion (L-PBF) heat transfer simulation, achieving sub-10-micrometer resolution for a 10 mm path in under 15 minutes on a single GPU. This makes a transformative step forward across all domains essential to engineering software.
@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 partially resolved the issue of labor-intensive programming in Software 1.0 and has significantly advanced the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies in the fields of engineering and science presents unique challenges. One such challenge is the relatively low generalization accuracy of machine learning (ML)-based solvers compared to traditional numerical methods [4, 5, 6, 7, 8]. After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9]. Another challenge is the curse of dimensionality that arises when solving inverse design problems. As the design parameter space grows, computation requirements easily become prohibitive. Furthermore, the complexity of physical simulations continues to increase. For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis within a reasonable simulation time [10, 11, 12].
@@
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibrating in science/engineering problems."
@@
We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0. The key ideas behind INNs are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function. Figure 1 shows a pictorial description of the approach. When the input domain is discretized with a regular mesh (see the special case in Figure 1), an INN leverages tensor decomposition (TD) [13, 14, 15, 16] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear. INN facilitates three major tasks in computational science and engineering: 1) trains on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solves high-dimensional parameterized partial differential equations (PDEs), and 3) calibrates uncertain parameters rapidly and accurately.
[…]
@@
Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Network (INN). The INN box illustrates the graph representation of the input domain discretized with an arbitrary Euclidean mesh (left) and a regular mesh as a special case (right). Node and edge features are given from the discretized mesh. After $Q$-step message passing, each node $j$ will store an interpolation function $N^{[Q](j)}(\bm{x})$. Finally, the readout operation sums the product of the interpolation functions and nodal values $\bm{u}^{(j)}$. The superscripts with square brackets [] and parentheses () denote the message passing step and graph node index, respectively. The interpolation operator $\mathcal{J}$ denotes that $\mathcal{J}\bm{u}(\bm{x})$ is a function that interpolates discrete values of $\bm{u}^{(j)}$. The INN trainer employs data-driven loss functions (e.g., mean squared error loss for regression) while the INN solver adopts a residual loss of a partial differential equation (PDE). A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
Figure 2: Benchmark comparison of the single-scale finite difference method (FDM) solver [18], the variational multiscale FEM solver [11], and the INN solver. The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed. A 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM. Detailed problem definition and explanation can be found in Section 3. All solvers are graphics processing units (GPU) optimized with the JAX library [19].
2 INN: Unification of Training/Learning, Solving, and Calibrating
@@
Interpolating Neural Network (INN) generalizes numerical methods such as the finite element method (FEM) [20, 21, 22] as a subset of deep neural network (DNN) via graph network operations. Although different groups of researchers have developed the two (FEM and DNN), they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibrating, as illustrated in Figure 1.
@@
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of an INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input – as illustrated in the INN box of Figure 1. The conversion from mesh to graph has been widely accepted in the literature [23, 24, 25, 26]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each segment occupies a subspace of the $I$-dimensional input domain. For instance in Figure 1, the Euclidean mesh consists of segments of hexahedrons and triangular pillars.
Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$ that stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
@@
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [27]. In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state. A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
@@
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information (SI) Section 1.1 for visual illustrations). It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$. The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$. This can be any interpolation method such as Lagrange polynomials [28, 20], meshfree [29, 30, 31], and splines [32, 33]. The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [34, 35] and convolution hierarchical deep learning neural networks [7, 15]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 15]. See SI Section 1 for various choices of the message passing operation.
@@
$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L},$ (2)
where the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$). The interpolation operator $\mathcal{J}$ will designate an interpolated field output throughout this paper. The readout operation can be written as a tensor contraction (or matrix multiplication):
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. Here, $\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second dimension of tensor $\bm{U}$ and the vector (or the first dimension of tensor) $\bm{\mathcal{X}(x)}$, while $\bm{U}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication. The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of the INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [21, 22]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating.
@@
When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [13]. One of the widely used TD methods is Tucker decomposition [36, 37]. It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$). The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
@@
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses data. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$. As illustrated in Figure 1’s special case box, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
When there are more than one outputs $(L>1)$, the interpolated field becomes a vector of $L$ elements:
@@
$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T},$
where $\mathcal{J}\prescript{}{i}{\bm{u}}_{l}\in\mathbb{R}^{M_{i}\times 1}$. The trainable parameters are the core tensors $\mathcal{G}_{l}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}},l=1,\dots,L$, and the nodal values $\prescript{}{i}{\bm{U}}_{l}\in\mathbb{R}^{M_{i}\times J_{i}},l=1,\dots,L$, yielding a total count of $L\left(\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}\right)$ that scales linearly with the nodal discretization $J_{i}$.
@@
Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [38, 39, 40] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements. If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the Tucker decomposition in Eq. 6 becomes:
$\mathcal{J}\bm{u}(\bm{x})=\sum^{M}_{m=1}{[\mathcal{J}\prescript{m}{1}{\bm{u}}(x_{1})\odot\dots\odot\mathcal{J}\prescript{m}{I}{\bm{u}}(x_{I})]},$ (7)
where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represents element-wise multiplication. In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is one-dimensional. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$. Considering the trainable parameters of multi-layer perception (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output $(L=1)$ relationship. The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization. We assume that the nodal coordinates $(\bm{x}^{(j)})$ are fixed.